show me how to do this equation:3x+5=20 and this one too ok:y/2+3=9
Answer by matty19925(1)

3x+5=20
x=5 because there is no sign between the 3 and the x so it is multiply, so your answer is...3x+5=20
3x+5-5=20-5
20-5=15
x=15